1.a) Graph a PPC by connecting the following points in order: Good A Good B: 0 5 10 14 15 0 1.b) What is the opportunity cost of the 7th unit of Good B? The 2nd?
Answer #1

1 a)

10 9876543210 p009

b) Opportunity cost of 7th unit of good B = (10 - 5) / (8 - 6) = 5 /2 = 2.5 units of good A

     Opportunity cost of 2nd unit of good B = (15 - 4) / (3 - 0) = 1 /3 = 0.33 units of good A